OpenStudy (anonymous):

Solve using linear systems. 2(x-1) - 3(y-3) = 0 3(x+2) - (y-7) = 20

OpenStudy (mertsj):

2x-2-3y+9=0 2x-3y=-7 3x+6-y+7=20 3x-y=7

OpenStudy (mertsj):

So now we have: 2x-3y= -7 3x - y = 7

OpenStudy (anonymous):

yeaa

OpenStudy (mertsj):

Let's take that second equation and solve it for y. We would get y=3x-7

OpenStudy (mertsj):

Are you with me so far?

OpenStudy (anonymous):

wouldn't it be -y = -3x + 7

OpenStudy (anonymous):

oh wait.. y = -(-3x+7) = what u said yes go on

OpenStudy (mertsj):

now multiply by -1 to get y

OpenStudy (mertsj):

So let's take that value for y which is 3x-7 and put it in the first equation in place of y

OpenStudy (mertsj):

2x-3y=-7 2x-3(3x-7)=-7

OpenStudy (anonymous):

everything is clear till now

OpenStudy (mertsj):

2x-9x+21=-7 -7x=-28 x=4 But y = 3x-7 y=3(4)-7 y=12-7 y=5 So the solution of the system is the ordered pair (4,5) and if you graphed the two equations you were given the lines would cross at (4,5)

OpenStudy (anonymous):

damn, tysm

OpenStudy (mertsj):

you are welcome